OpenStudy (anonymous):

An equation and the steps to solve it are shown below: 3x - alt= (8x - 2) = 4 Step 1: 3x - 4x + 1 = 4 Step 2: x + 1 = 4 Step 3: x + 1 - 1 = 4 - 1 Step 4: x = 3 Which step was the error made and why? (1 point) Question 19 options: 1) Step 1, because incorrect signs have been used 2) Step 2, because the terms have been added incorrectly 3) Step 3, because numbers have been added to both sides of the equation 4) No error was made

OpenStudy (anonymous):

what does - alt mean?

OpenStudy (anonymous):

@bobmath sorry its supposed to be 1 - 2

OpenStudy (anonymous):

ok so 3x - 1 = (8x - 2) - 4 ?

OpenStudy (anonymous):

or + 4 because I'm confused by the equals sign before the 4

OpenStudy (anonymous):

3x - 1= (8x - 2) = 4 @bobmath - 2

OpenStudy (anonymous):

Step 1 is wrong. If 8x - 2 = 4, then 3x - 4x + 1 cannot equal 4

OpenStudy (anonymous):

Thank You!

OpenStudy (anonymous):

you're welcome